# Blessing of Multilinguality: A Systematic Analysis of Multilingual In-Context Learning

### v1

@@ Abstract
While multilingual large language models generally perform adequately, and sometimes even rival English performance on high-resource languages (HRLs), they often significantly underperform on low-resource languages (LRLs; Costa-jussà et al., 2022). Among several prompting strategies aiming at bridging the gap, multilingual in-context learning (ICL; Shi et al., 2023) has been particularly effective when demonstration in target languages is unavailable. However, there lacks a systematic understanding when and why it works well.
In this work, we systematically analyze multilingual ICL, using demonstrations in HRLs to enhance cross-lingual transfer. We show that demonstrations in mixed HRLs consistently outperform English-only ones across the board, particularly for tasks written in LRLs. Surprisingly, our ablation study show that the presence of irrelevant non-English sentences in the prompt yields measurable gains, suggesting the effectiveness of multilingual exposure itself. Our results highlight the potential of strategically leveraging multilingual resources to bridge the performance gap for underrepresented languages.
@@
Yilei Tu Andrew Xue Freda Shi David R. Cheriton School of Computer Science, University of Waterloo 200 University Ave W, Waterloo, ON N2L 5Z5, Canada {yileitu,a8xue,fhs}@uwaterloo.ca
@@ 1 Introduction
Prior work has introduced two common ICL strategies for non-English languages: (i) translating the target question into English and performing English-only ICL (Shi et al., 2023; Ahuja et al., 2023; Razumovskaia et al., 2024, inter alia), and (ii) providing demonstrations in the target language (in-language demonstrations; Fu et al., 2022; Qin et al., 2023; Huang et al., 2023; Zhang et al., 2024b). Both strategies have critical shortcomings: (i) may suffer from information loss in translation due to nuanced language gap (Zhang et al., 2023; Poelman and de Lhoneux, 2024) or the unavailability of high-quality translation systems for extremely low-resource languages (LRLs), whereas (ii) may become infeasible due to data scarcity in LRLs. As alternatives, when presenting LLMs with the problems in the target language, English demonstrations (Fig. 1(a)) lead to poor performance on LRLs, whereas demonstrations in multiple high-resource languages (HRLs; Fig. 1(b)) can be more effective, even when there are little alphabetical overlap between the demonstration and target languages (Shi et al., 2023); however, the underlying reasons on why it works remain unclear.
In this work, we systematically analyze multilingual ICL through a set of controlled experiments. Each test question is paired with a set of semantically equivalent demonstrations, while the demonstration languages vary according to different ICL modes (Tab. 1). We compare the performance differences across four ICL modes (§§ 4.1 and 4.2): English, individual-HRL, multilingual (i.e., mixed HRLs), and in-language demonstrations (Fig. 2). To disentangle the impact of demonstration language from other confounding factors (e.g., interactions between demonstration languages and in-domain demonstrations), we conduct additional control experiments by adding irrelevant sentences in various languages into English-only in-domain demonstrations (§ 4.3). We find that
In-context demonstrations in HRLs, especially in languages with non-Latin writing system such as Chinese and Japanese, can more effectively transfer knowledge compared to English ICL mode, leading to performance improvement on answering questions in all languages, especially in LRLs. This finding is generalizable enough across different domains and various LLMs.
Demonstrations in mixed HRLs is more robust and effective compared to that in a single HRL, in terms of average accuracy boosting on different tasks. This strategy is favored.
@@ 2 Background: Multilingual ICL
In this section, we review the basic approaches of ICL with instruction-tuned LLMs (§ 2.1) and introduce the multilingual prompting modes (§ 2.2) that we evaluate in this work.
@@ 2.2 Multilingual Prompting Modes
The ground-truth labels in quantitative LLM benchmarks (Ponti et al., 2020; Cobbe et al., 2021, inter alia) are typically language-agnostic (such as numbers) or represented in a single word (such as Yes/No). In such cases, the verbalizer is an identity. For answers requiring reasoning steps, is CoT in English, as there has been strong evidence that MLLMs performs better when generating English (Shi et al., 2023; Qin et al., 2023; Huang et al., 2023, inter alia). For the same reasons, is always presented in English as well.
@@
The English mode. The demonstrations are always in English (Figs. 1(a) and 2a).
The Monolingual mode(s). The demonstrations are always presented in a single non-English high-resource language such as Chinese (Fig. 2b).
The Multilingual mode. From a predefined list of high-resources languages , languages are randomly selected, which, together with the sampled indices, determine the contents and languages of the demonstrations (Figs. 1(b) and 2c).
The Native mode. The demonstrations are in the same language as the test question (Fig. 2d).
@@ 3 Experiment Setups
Models. We evaluate state-of-the-art instruction-tuned LLMs with about 8 billion parameters, which have officially claimed multilingual capabilities in model release: Llama3-8B-Instruct, Llama3.1-8B-Instruct (Dubey et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024a), Qwen2.5-7B-Instruct (Yang et al., 2024b), Mistral-NeMo-12B-Instruct (MistralAI, 2024) and Aya-Expanse-8B (Dang et al., 2024). For additional references, we evaluate OpenAI closed-sourced commercial models, including GPT3.5-turbo (OpenAI, 2022) and GPT4o-mini (OpenAI, 2024). Detailed model cards can be found in § A.1.
@@
[…]
| See Fig. 1 for examples. | ||||||
[…]
@@
Datasets. We evaluate the models using multilingual benchmarks from three distinct domains: (1) MGSM Shi et al. (2023), a benchmark of 250 grade-school math problems sampled from the English GSM8K (Cobbe et al., 2021) and translated into 10 additional languages by expert native speakers. (2) XCOPA Ponti et al. (2020), a commonsense reasoning benchmark that extends the COPA dataset (Roemmele et al., 2011) to additional languages.111In this work, we merge the English COPA and XCOPA datasets together, which we still refer to as XCOPA. (3) XL-WiC Raganato et al. (2020), a cross-lingual word-in-context understanding dataset spanning languages, where models are expected to tell whether a polysemous word retains the same meaning in two contexts.
@@
MGSM and XCOPA are parallel where each corresponding datapoint across different language splits contains semantically equivalent content, allowing us to minimize semantic confounders in our experimental design. XL-WiC is language-specific and translation-variant thus naturally non-parallel. Dataset properties and examples are in Tabs. 1 and 6. Details of data curation are in § A.2.
@@
Languages. Languages with large-scale digitized data resources on web are known as high-resource languages (HRLs; Bender, 2019), which are exemplified by English, Spanish, Chinese, among others. In contrast, low-resource languages (LRL) have scarce accessible data Costa-jussà et al. (2022). However, a universal standard for dichotomizing languages as either high- or low-resource has not been set (Bender, 2019; Joshi et al., 2020; Hedderich et al., 2021). Moreover, none of the models we evaluate has disclosed the language distributions in their training corpora. As a workaround, we define our preset HRL list as the union of the 20 most frequent languages in Llama2 Touvron et al. (2023) and PaLM Chowdhery et al. (2023), and classify languages out of the HRL list as LRLs. Details of preset HRL list can be found in § A.3.
@@
Prompts. Following Shi et al. (2023), we use examples for demonstration for any test question. For each multilingual dataset (Tab. 1), we first sample index lists of length all at once, where the index range is We allocate the -th index list to the set of test questions for the same index across all language splits. The training-set index list, together with the specified languages,222 For the Multilingual mode, we apply the same sampling procedure to generate HRL code lists of length , drawing from the available HRLs in the dataset. jointly determines the content and language of the demonstration for each testing example (Fig. 2). This approach both ensures linguistic diversity for multilingual prompting and, whenever applicable, mitigates confounding factors that come with semantic inconsistency across examples. All interface words (see Tab. 1) are in the same language as the examples rather than in English.
@@
Inference and metrics. Throughout this work, we use greedy decoding for inference, selecting the token with the highest probability at each step. For MGSM in need of CoT, we set the maximum token length to 500; for XCOPA and XL-WiC, we set it to 10, as we expect the answers to be short. We use exact match accuracy as our evaluation metric: for MGSM, we extract the last numeral in the response. For XCOPA and XL-WiC, we extract label (expected output) in the response (Tab. 1).
@@ 4.1 Multilingual Prompts Surpass English
Results. We first compare the -shot performance with English, Multilingual and Native ICL modes on open-sourced MLLMs and commercial OpenAI models (Fig. 3), with Tab. 2 presenting the detailed performance of three selected MLLMs across various LRLs. Overall, Multilingual mode outperforms English mode, both for individual LRLs and on average. In out of cases (Fig. 3), Multilingual mode achieves higher accuracy than English one. This phenomenon is evident even for GPT4o-mini, one of the currently strongest LLMs Chiang et al. (2024), and for HRLs as well (see Tabs. 11, 12 and 13 in § B.1 for HRL accuracies). Extending the results of Shi et al. (2023) that Multilingual mode generally outperforms English mode for PaLM (Chowdhery et al., 2023) and Codex (Chen et al., 2021), our results confirm this trend is a general phenomenon across various MLLMs and datasets.
Outstanding performance of the Native mode and the practical unfeasibility. Admittedly, in out of comparisons, Native mode outperforms Multilingual mode (Fig. 3), which aligns with the machine-learning intuition that in-domain data, in terms of both genre and language, are more promising for model performance (Liu et al., 2024). However, domain-specific datasets for LRLs are often difficult to obtain due to the scarcity of native speakers or professional translators (Costa-jussà et al., 2022; NLLB-Team, 2024); therefore, in practice, it is usually challenging to provide high-quality demonstrations in the same language and domain as the test question. In contrast, annotations makes the HRL-Multilingual mode more feasible in many scenarios.
Hypothesis Tests. To verify whether the improvement is statistically significant, we conduct McNemar’s test McNemar (1947)—the null hypothesis means no significant accuracy difference between the baseline (English) and the compared mode. Let denote the number of cases where the baseline is correct while the compared mode is incorrect, denotte the number of cases where the baseline mode is incorrect while the compared mode is correct. We calculate the corrected version Edwards (1948) of the McNemar’s statistic:
which has a chi-squared distribution with one degree of freedom. Significant -test results provide strong evidence to reject the null hypothesis of no accuracy improvement. Our results results (Tab. 2 and Tabs. 14, 15 and 16 in § B.1) indicate that both Multilingual and Native modes significantly outperform the English mode.
@@ 4.2 Ablation Study: Non-English Monolingual Prompts Are Effective
With the success of the Multilingual mode, we investigate whether the improvement comes from the introduction of multiple languages or simply from a single non-English HRL. Specifically, we compare several Monolingual modes including Chinese (for all three datasets), French, Japanese (both for MGSM and XL-WiC) and Italian (for XCOPA). We select French and Italian because they are both European languages and thus share considerable subword overlap with English; in contrast, Chinese and Japanese exhibit little subword overlap with Latin-script languages, but there is a substantial overlap between them two due to their shared use of Chinese characters (or kanji). This language selection allows us to analyze simultaneously the impact of the writing system (or subword overlap) on ICL performance.
Results. All HRL-Monolingual modes outperform English in a considerable number of comparisons (Fig. 4). This finding also holds for HRL evaluations (see Tabs. 11, 12 and 13 in § B.1 for raw accuracies of each language). Among all Monolingual modes, Chinese performs the best, matching Multilingual mode with 17 out of 21 comparisons when accuracy surpasses that of English. Japanese also frequently outperforms English. Extending the findings of Turc et al. (2021) that non-English languages are more effective than English in pretraining and fine-tuning based cross-lingual transfer, our results suggest that non-English languages, particularly those with non-Latin scripts, may be more effective under the prompting scheme as well.
However, the Multilingual mode exhibits stronger robustness, outperforming Chinese in 14 out of 21 LRL cases. Same trends applies to HRLs. The results of the hypothesis test further confirm the robustness: for both LRL and HRL splits, the Multilingual mode exhibits the highest number of significant results relative to other Monolingual modes (Tabs. 14, 15 and 16 in § B.1). Intuitively, we hypothesize that Multilingual mode functions like an “average” of individual HRL-Monolingual modes, making it most robust, and thus it outperforms English most frequently and achieves the highest overall average accuracy.
Following Tang et al. (2024), we further identify ICL-mode-specific neurons and find that the neurons activated by Multilingual overlap most with those activated by Native among other modes. This further explains why Multilingual could achieve performance comparable to Native. See App. C for details.

### v3

@@ Abstract
While multilingual large language models generally perform adequately, and sometimes even rival English performance on high-resource languages (HRLs), they often significantly underperform on low-resource languages (LRLs; Costa-jussà et al., 2022). Among several prompting strategies aiming at bridging the gap, multilingual in-context learning (ICL; Shi et al., 2023) has been particularly effective when demonstration in target languages is unavailable. However, there lacks a systematic understanding of when and why it works well.
In this work, we systematically analyze multilingual ICL, using demonstrations in HRLs to enhance cross-lingual transfer. We show that demonstrations in mixed HRLs consistently outperform English-only ones across the board, particularly for tasks written in LRLs. Surprisingly, our ablation study shows that the presence of irrelevant non-English sentences in the prompt yields measurable gains, suggesting the effectiveness of multilingual exposure itself. Our results highlight the potential of strategically leveraging multilingual resources to bridge the performance gap for underrepresented languages.
@@
Yilei Tu1 Andrew Xue1 Freda Shi1,2,3 1David R. Cheriton School of Computer Science, University of Waterloo 2Vector Institute 3Canada CIFAR AI Chair yileitu.tt@gmail.com, fhs@uwaterloo.ca
@@ 1 Introduction
Prior work has introduced two common ICL strategies for non-English languages: (i) translating the target question into English and performing English-only ICL (Shi et al., 2023; Ahuja et al., 2023; Razumovskaia et al., 2024, inter alia), and (ii) providing demonstrations in the target language (in-language demonstrations; Fu et al., 2022; Qin et al., 2023; Huang et al., 2023; Zhang et al., 2024b). Both strategies have critical shortcomings: (i) may suffer from information loss in translation due to nuanced language gap (Zhang et al., 2023; Poelman and de Lhoneux, 2024) or the unavailability of high-quality translation systems for extremely low-resource languages (LRLs), whereas (ii) may become infeasible due to data scarcity in LRLs. As alternatives, when presenting LLMs with problems in the target language, English demonstrations (Fig.˜1(a)) lead to poor performance on LRLs, whereas demonstrations in multiple high-resource languages (HRLs; Fig.˜1(b)) can be more effective, even when there is little alphabetical overlap between the demonstration and target languages (Shi et al., 2023); however, the underlying reasons on why it works remain unclear.
In this work, we systematically analyze multilingual ICL through a set of controlled experiments. Each test question is paired with a set of semantically equivalent demonstrations, while the demonstration languages vary according to different ICL modes (Tab.˜1). We compare the performance differences across four ICL modes (§§˜4.1 and 4.2): English, individual-HRL, multilingual (i.e., mixed HRLs), and in-language demonstrations (Fig.˜2). To disentangle the impact of demonstration language from other confounding factors (e.g., interactions between demonstration languages and in-domain demonstrations), we conduct additional control experiments by adding irrelevant sentences in various languages into English-only in-domain demonstrations (§˜4.3). We find that
In-context demonstrations in HRLs, especially in languages with non-Latin writing systems such as Chinese and Japanese, can more effectively transfer knowledge compared to English ICL mode, leading to performance improvement in answering questions in all languages, especially in LRLs. This finding is generalizable enough across different domains and various LLMs.
Demonstrations in mixed HRLs are more robust and effective compared to those in a single HRL, in terms of average accuracy boosting on different tasks. This strategy is favored.
@@ 2 Background: Multilingual ICL
In this section, we review the basic approaches of ICL with instruction-tuned LLMs (§˜2.1) and introduce the multilingual prompting modes (§˜2.2) that we evaluate in this work.
@@ 2.2 Multilingual Prompting Modes
The ground-truth labels in quantitative LLM benchmarks (Ponti et al., 2020; Cobbe et al., 2021, inter alia) are typically language-agnostic (such as numbers) or represented in a single word (such as Yes/No). In such cases, the verbalizer is an identity. For answers requiring reasoning steps, is CoT in English, as there has been strong evidence that MLLMs perform better when generating English (Shi et al., 2023; Qin et al., 2023; Huang et al., 2023, inter alia). For the same reasons, is always presented in English as well.
@@
The English mode. The demonstrations are always in English (Figs.˜1(a) and 2a).
The Monolingual mode(s). The demonstrations are always presented in a single non-English high-resource language such as Chinese (Fig.˜2b).
The Multilingual mode. From a predefined list of high-resources languages , languages are randomly selected, which, together with the sampled indices, determine the contents and languages of the demonstrations (Figs.˜1(b) and 2c).
The Native mode. The demonstrations are in the same language as the test question (Fig.˜2d).
@@ 3 Experiment Setups
Models. We evaluate state-of-the-art instruction-tuned LLMs with about 8 billion parameters, which have officially claimed multilingual capabilities in model release: Llama3-8B-Instruct, Llama3.1-8B-Instruct (Dubey et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024a), Qwen2.5-7B-Instruct (Yang et al., 2024b), Mistral-NeMo-12B-Instruct (MistralAI, 2024) and Aya-Expanse-8B (Dang et al., 2024). For additional references, we evaluate OpenAI closed-sourced commercial models, including GPT3.5-turbo (OpenAI, 2022) and GPT4o-mini (OpenAI, 2024). Detailed model cards can be found in §˜A.1.
@@
| MGSM | Mathematical Reasoning | Numerals | ||||
| See Fig.˜1 for examples. | ||||||
| XCOPA | Commonsense Reasoning | “1” / “2” | ||||
@@
| XQuAD | Extractive QA | Text | ||||
| Passage: The Panthers defense gave up just 308 points, ranking sixth in the league, | ||||||
| Question: How many points did the Panthers defense surrender? Answer: 308. | ||||||
@@
Datasets. We evaluate the models using multilingual benchmarks from distinct domains: (1) MGSM Shi et al. (2023), a benchmark of grade-school math problems sampled from the English GSM8K (Cobbe et al., 2021) and translated into 10 additional languages by expert native speakers. (2) XCOPA Ponti et al. (2020), a commonsense reasoning benchmark that extends the COPA dataset in English (Roemmele et al., 2011) to additional languages. (3) XL-WiC Raganato et al. (2020), a cross-lingual word-in-context understanding dataset spanning languages, where models are expected to tell whether a polysemous word retains the same meaning in two contexts. (4) XQuAD Artetxe et al. (2020) is designed to evaluate cross-lingual question answering performance, based on the English SQuAD Rajpurkar et al. (2016) dataset and professionally translated into 10 languages.
@@
MGSM, XCOPA and XQuAD are parallel where each corresponding datapoint across different language splits contains semantically equivalent content, allowing us to minimize semantic confounders in our experimental design. XL-WiC is language-specific and translation-variant thus naturally non-parallel. Dataset properties and examples are in Tabs.˜1 and 6. Details of data curation are in §˜A.2.
@@
Languages. Languages with large-scale digitized data resources on the web are known as high-resource languages (HRLs; Bender, 2019), which are exemplified by English, Spanish, and Chinese, among others. In contrast, low-resource languages (LRL) have scarce accessible data Costa-jussà et al. (2022). However, a universal standard for dichotomizing languages as either high- or low-resource has not been set (Bender, 2019; Joshi et al., 2020; Hedderich et al., 2021). Moreover, none of the models we evaluate has disclosed the language distributions in their training corpora. As a workaround, we define our preset HRL list as the union of the 20 most frequent languages in Llama2 Touvron et al. (2023) and PaLM Chowdhery et al. (2023), and classify languages out of the HRL list as LRLs. Details of the preset HRL list can be found in §˜A.3.
@@
Prompts. Following Shi et al. (2023), we use examples for demonstration for any test question. For each multilingual dataset (Tab.˜1), we first sample index lists of length all at once, where the index range is We allocate the -th index list to the set of test questions for the same index across all language splits. The training-set index list, together with the specified languages,111For the Multilingual mode, we apply the same sampling procedure to generate HRL code lists of length , drawing from the available HRLs in the dataset. jointly determines the content and language of the demonstration for each testing example (Fig.˜2). This approach both ensures linguistic diversity for multilingual prompting and, whenever applicable, mitigates confounding factors that come with semantic inconsistency across examples. All interface words (see Tab.˜1) are in the same language as the examples rather than in English.
@@
Inference and metrics. Throughout this work, we use greedy decoding for inference, selecting the token with the highest probability at each step. For MGSM in need of CoT, we set the maximum token length to ; for XCOPA and XL-WiC, we set it to 10, as we expect the answers to be short; for XQuAD, it is . We use exact match accuracy as our evaluation metric: for MGSM, we extract the last numeral in the response; for XCOPA and XL-WiC, we extract the label (expected output) in the response (Tab.˜1); for XQuAD, a response of an MLLM is considered correct if it contains the gold answer as a substring after normalization.
@@ 4.1 Multilingual Prompts Surpass English
Results. We first compare the -shot performance with English, Multilingual and Native ICL modes on open-sourced MLLMs and commercial OpenAI models (Fig.˜3), with Tab.˜2 presenting the detailed performance of three selected MLLMs across various LRLs. Overall, Multilingual mode outperforms English mode, both for individual LRLs and on average. In out of cases (Fig.˜3), Multilingual mode achieves higher accuracy than English one. This phenomenon is evident even for GPT4o-mini, one of the currently strongest LLMs Chiang et al. (2024), and for HRLs as well (see Tabs.˜11, 12 and 13 in §˜B.1 for HRL accuracies). Extending the results of Shi et al. (2023) that Multilingual mode generally outperforms English mode for PaLM (Chowdhery et al., 2023) and Codex (Chen et al., 2021), our results confirm this trend is a general phenomenon across various MLLMs and datasets.
Outstanding performance of the Native mode and the practical unfeasibility. Admittedly, in out of comparisons, Native mode outperforms Multilingual mode (Fig.˜3), which aligns with the machine-learning intuition that in-domain data, in terms of both genre and language, are more promising for model performance (Liu et al., 2024). However, domain-specific datasets for LRLs are often difficult to obtain due to the scarcity of native speakers or professional translators (Costa-jussà et al., 2022; NLLB-Team, 2024); therefore, in practice, it is usually challenging to provide high-quality demonstrations in the same language and domain as the test question. In contrast, annotations make the HRL-Multilingual mode more feasible in many scenarios.
Hypothesis Tests. To verify whether the improvement is statistically significant, we conduct McNemar’s test McNemar (1947)—the null hypothesis means no significant accuracy difference between the baseline (English) and the compared mode. Let denote the number of cases where the baseline is correct while the compared mode is incorrect, and denote the number of cases where the baseline mode is erroneous while the compared mode is correct. We calculate the corrected version Edwards (1948) of the McNemar’s statistic:
which has a chi-squared distribution with one degree of freedom. Significant -test results provide strong evidence to reject the null hypothesis of no accuracy improvement. Our results results (Tab.˜2 and Tabs.˜15, 16, 17 and 18 in §˜B.1) indicate that both Multilingual and Native modes significantly outperform the English mode.
@@ 4.2 Ablation Study: Non-English Monolingual Prompts Are Effective
With the success of the Multilingual mode, we investigate whether the improvement comes from the introduction of multiple languages or simply from a single non-English HRL. Specifically, we compare several Monolingual modes including Chinese (for all datasets), French, Japanese (both for MGSM and XL-WiC), Italian (for XCOPA) and German (for XQuAD). We select French, Italian and German because they are European languages and thus share considerable subword overlap with English; in contrast, Chinese and Japanese exhibit little subword overlap with Latin-script languages, but there is a substantial overlap between the two due to their shared use of Chinese characters (or Hanzi, Kanji). This language selection allows us to analyze simultaneously the impact of the writing system (or subword overlap) on ICL performance.
Results. All HRL-Monolingual modes outperform English in a considerable number of comparisons (Fig.˜4). This finding also holds for HRL evaluations (see Tabs.˜11, 12, 13 and 14 in §˜B.1 for raw accuracies of each language). Among all Monolingual modes, Chinese performs the best, with out of comparisons when accuracy surpasses that of English. Japanese also frequently outperforms English. Extending the findings of Turc et al. (2021) that non-English languages are more effective than English in pretraining and fine-tuning based cross-lingual transfer, our results suggest that non-English languages, particularly those with non-Latin scripts, may be more effective under the prompting scheme as well.
However, the Multilingual mode exhibits stronger robustness, outperforming Chinese in out of LRL cases. The same trend applies to HRLs. The results of the hypothesis test further confirm the robustness: for both LRL and HRL splits, the Multilingual mode exhibits the highest number of significant results relative to other Monolingual modes (Tabs.˜15, 16, 17 and 18 in §˜B.1). Intuitively, we hypothesize that Multilingual mode functions like an “average” of individual HRL-Monolingual modes, making it most robust, and thus it outperforms English most frequently and achieves the highest overall average accuracy.
Following Tang et al. (2024), we further identify ICL-mode-specific neurons and find that the neurons activated by Multilingual overlap most with those activated by Native among other modes. This further explains why Multilingual could achieve performance comparable to Native. See App.˜C for details.